OpenStudy (anonymous):

CALCULUS : NORMAL CURVE Suppose \(f(x)= -\frac{ 1 }{ 4 }x^2+3x+4\). a. i. Write down \(f'(x)\) in simplest form. ii. Find the equation of the normal to \(f\) at \((2,9)\). iii. If this normal cuts \(f\) at another point A, find the coordinates of A.

OpenStudy (mathmale):

Nice that you've used Equation Editor. Neat, accurate results! What's the derivative of 4? What's the derivative of 3x?

OpenStudy (anonymous):

Hold on; my question is not finished. OS is so slow. -_-

OpenStudy (anonymous):

a. i. \(f'(x)=-\frac{ 1 }{ 2 }x+3\)

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Yeah, power rule 

OpenStudy (anonymous):

All right. So for ii., how should I start?

jimthompson5910 (jim_thompson5910):

ii) step 1) plug in x = 2 into the f ' (x) equation. Evaluate to find the value of the slope of the tangent line. Call it `m` step 2) Compute `-1/m` to find the perpendicular slope Step 3) use either point slope or slope intercept form to find the equation of the normal line

OpenStudy (anonymous):

\(f'(2)=2 \) therefore m= -1/2

OpenStudy (anonymous):

y-9=-1/2(x-2) or y=-1/2+10

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

the second equation I think you meant to say `y = (-1/2)x+10`

jimthompson5910 (jim_thompson5910):

if so, then they both look good. I would go with \(\Large y = -\frac{1}{2}x+10\)

OpenStudy (anonymous):

yes, i forgot the x :)

OpenStudy (anonymous):

then for iii., do i just equate the f(x) and f'(x)?

OpenStudy (anonymous):

if so, i got (12, 4) as coordinates

jimthompson5910 (jim_thompson5910):

you have the system of equations with f and the normal line so these two equations y = -1/4x^2+3x+4 y = -1/2x + 10 solve this system. One of the points will be (2,9). They want you to find the other point

OpenStudy (anonymous):

oh i meant the normal curve and the function not the derivative

OpenStudy (anonymous):

i'll show my work

jimthompson5910 (jim_thompson5910):

(12,4) is correct. That is the other point

OpenStudy (anonymous):

okay thanks! :)

jimthompson5910 (jim_thompson5910):

no problem